OpenStudy (anonymous):

find three consecutive odd integers such that 3 times the sum of the first two integers is equal to 4 times the thrid integer

OpenStudy (anonymous):

yes odd

OpenStudy (anonymous):

now what

OpenStudy (mertsj):

x=first integer x+2 = second consecutive odd integer x+4 = third consecutive odd integer 3 times the sum of the first two...write that

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

how do i put them together

OpenStudy (anonymous):

@Mertsj can you help please

OpenStudy (mertsj):

Did you write what I said?

OpenStudy (anonymous):

yes

OpenStudy (mertsj):

Post it so I can see it.

OpenStudy (anonymous):

x=1st integer x+2 = 2nd integer x+4= 3rd integer x+x+2=2x+3 than i multiply that by 3 and get 6x+6

OpenStudy (mertsj):